RunPlanDaily CREX

From PREX Wiki
Jump to: navigation, search



CREX Main << Run Coordinator Tools

Global Run Plan

Current Daily Run Plans

The current day is indicated by a color.

Day, Date Shift Tasks (Responsible Party) Comments/Status
Monday, Jan 6 Owl
Day
Swing Moller Lock, FFB low current, Optics Raster Iy issue resolved, all same as Dec.
Tuesday, Jan 7 Owl Optics Larm det ul base replaced
Day Mott, Spin Dance Vac gauge readout 1-e12, vac adjacent fine (not urgent, to be fix later)
Swing Spin Dance, 150uA to dump Experts: (Simona, Eric) , (Caryn, Devi)
Wednesday, Jan 8 Owl 150uA, bcm/det calibration, damping measure, bmw initialization...production Experts: Caryn, Devi, Ye
Day Hall Access, fix compton T controller. OPS down:RF recovery, NL RF cav gradient, energy loss, broken klystron, failed arc det.
Swing Tune through chicane attempt, bleed-through from B discovered, returned to straight through end of shift, but compton rates too high Experts: Compton(Dave, Adam, Amali), BMW(Ye,Victoria,Caryn)
Thursday, Jan 9 Owl Low current BMW tune 1H01 test Couldn't send beam due to compton rates
Day halld laser inj access, bleedthrough B into A injector work prebuncher, chopper.
Swing compton tuning through chicane...but match into hall was wrong. Couldn't send beam due to spot size limitations.
Friday, Jan 10 Owl Tuning match into Hall
Day "Production", injector adjustment, injector match
Swing injector match, hall match, hall on design, lower compton rates straight through, compton chicane tuning attempt ran out of time
Sat Jan 11 Owl Production, FFB gain issues (hi search limit), yield det decrease and width increase over night
Day detector alignment check and left dipole current adjustment. Production. LQ1 trip, recycle.
Swing 7:53PM HA to power permit, beam -> C
Sun Jan 12 Owl (figure out HC feedback again)
Day Radcon survey, Compton Laser work (TO HC feedback )
Swing Compton Laser work (TO HC feedback )
Mon Jan 13 Owl (TO HC feedback )
Day Vacuum group work ion pump, Compton laser work (TO HC feedback )
Swing Compton laser work, Hall Lockup, tell OPS PQB checkout in INJ when HA coming back online (TO HC feedback )
Tues Jan 14 Owl Production,
Day Compton chicane tuning, commissioning, Production Don's Run Plan
Swing BMW fine tune, Production Don's Run Plan
Wed Jan 15 Owl Production,
Day Compton chicane tuning, commissioning, Production Don's Run Plan
Swing BMW fine tune, Production Don's Run Plan
Th Jan 16 Owl Production,
Day Hall C tuning, production
Swing Production,
Fr Jan 17 Owl Compton setup, Production,
Day Machine tuning; return to production
Swing Production
Sat Jan 18 Owl Production,
Day Production
Swing Production
Sun Jan 19 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access
Swing Down due to target issue, controlled access
Mon Jan 20 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access
Swing Down due to target issue, controlled access
Tue Jan 21 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access
Swing Down due to target issue, controlled access
Wed Jan 22 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access Ca40, Ca48 target reinstalled
Swing
Thu Jan 23 Owl
Day Target cooldown
Swing Beam restoration, target commissioning, detector alignment check, Qsq measurement
Fri Jan 24 Owl BMW check, target thermal commissioning
Day Production!
Swing Production!
Sat Jan 25 Owl
Day
Swing
Sun Jan 26 Owl
Day Moller measurement
Swing Moller measurement

Previous Daily Run Plans

Instructions for RC

  • Let the day shift know what to do while you attend the MCC meeting
  • Copy the daily run plans for the previous week to the Run plan history CREX
  • Copy the Run Plan Template table to this page and change the dates
  • Update the plan above (if necessary) after the MCC meeting